pick up 2 piles of cards and put 1 in the middle of the pack and do it again and again repeat......
Try 5/16 x 4/15 x 3/14 = 1/56 = 0.010178571
What is the formula in probability?
Probability equals favorable outcomes divided by total number of outcomes.
No. of apples picked = 1
No. of red apples = 6
No. of yellow apples = 8
Total no. of apples = 14
Probability that the apple picked is red = No. of red / Total no. of red apples
= 6/14
= 3/7
= 0.42857142857142857142857142857143
Two dice a tossed find the probability that the numbers are the same given that at least one is 3?
There are two different ways to look at this that give two different answers. The first way is this. We already know that one of the dice is a three. So the question of whether the two dice show the same number becomes "Is the second die a 3?" And the probability of the second die being a 3 is 1/6. Without really making it clear, this method ASSUMES that it is the FIRST die that is constrained to be a 3. The other way is this. The question constrains AT LEAST ONE of the two die to be a 3, but we don't know which one. So, what you have to do is draw a chart, listing all the possibilities (1-6) for die #1 on the left and all the possibilities for die #2 (also 1-6)across the top. There are 36 possible outcomes. But in only 11 of them does EITHER of the two dice show a 3. And, among those 11, in only 1 does BOTH dice show a 3. Therefore, the probability is 1/11.
The answer depends on what is the probability of WHAT!
expected profit should be : 0.7($36,000) - 0.3($6,000) = $23,400
One card is made from a deck of 52 cards What is the probability that a king will be drawn?
The answer is 4/52 which can be simplified to 1/13.
Where is probability used in the real world?
Probability is used everywhere: Betting odds. Medical odds, (chance of survival or chance of side effect happening). Anywhere we calculate risks (insurances calculate premiums based on probability). Communication Networks
What is the probability of an event occurring to the probability that will not occur?
If the event in question is A and the probability of A occurring is P(A), then the probability of A not occurring is P(A')=1-P(A).
Real life application of quadratic equation?
How about the path a baseball takes when hit by a bat...
one of the many google easter eggs is doing a barrel roll if you search do a barrel roll on google the screen will do a 360 degree somersault
How do you find the standard deviation?
The standard deviation (?, pronounced sigma) of a set of values is a measure of how much the set of values deviates from the average of the values. To calculate ? of a complete set of values (as opposed to a sampling),...
How do you calculate probability?
Probability refers to the likelihood of an event occurring. As such, calculating the same involves dividing the chances of an event occurring by the probable number of times that it can occur.
Probability is a measure of the likelihood of an event. There are different ways to calculate probability it all depend on what probability you are trying to calculate. The general formula to calculate a probability is to divide the number of event you are trying to calculate the probability for by the total number of out comes.
Think of a dice the probability of rolling a 1 is equal to 1 (there is one way we can roll a one) divide by 6 ( the total number of possible out comes, i.e. 1/6.
The question is too broad. Please re-ask the question with more specifics.In general, you divide the number of anticipated outcomes by the number of possible outcomes. For instance, the probability of drawing an ace in a standard deck is 4 in 52, or 1 in 13, or 0.0769
What is the theoretical probability of getting 4 tails out of ten throws?
theoretical probability is one half experimental probability is four tenths this is because to find theoretical probability you need to do number of outcomes you were looking for over the number of outcomes possible experimental probability is number of turns that were what you were looking for over the number of turns
ok let's try this!
on any given date, probability that there is a student with a birthday is 30/365, so prob. that two students have that birthday is (30/365)^2. Tell me if I'm wrong, somebody!
So the problem that this does NOT happen is 1-(30/365)^2 which is 0.9932445...
I'm not sure about this next bit:
there are 365 days in which this probably won't occur, does that make a difference?
How do i find the percentage loss?
18.55% is overall loss Loss per annum compounding is 5.263% The fact that 4 years is given would indicate that the answer required is the annual compounding figure anmd not simply the first answer above. Overall loss divide the difference between the starting and closing figures by the starting figure Loss per annum use the formula S=P(1=i)power of n Where S is closing figure P is starting figure i is interest rate n is number of years I solved latter using a formula in excel within (fx)
I think it depends on how many students are thinking of alphabet letters at the same time :) The first student will think of a letter from the English alphabet. Then, the second student will think of a letter. The probability that the second student will select what the first student selected is 1/26.
It's still 1 in 6. It's always 1 in 6. Each roll is independent.
The chance is 1/6 as always.